Construct the confidence interval for the population mean mu. cequals0.95​, x overbar equals 15.2​, sigmaequals4.0​, and nequals55 A 95​% confidence interval for mu is
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 15.2

Population standard deviation =   \sigma = 4.0

Sample size = n =55

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96


Margin of error = E = Z\alpha/2 * ( \sigma /\sqrtn)

= 1.96 * ( 4.0 /  \sqrt55 )

= 1.0571
At 95% confidence interval estimate of the population mean
is,

\bar x - E < \mu < \bar x + E

15.2 - 1.0571 <  \mu < 15.2+ 1.0571

14.1429 <  \mu < 16.2571

( 14.1429 , 16.2571)